# seg2med: a segmentation-based medical image generation framework using denoising diffusion probabilistic models

### v1

@@ Abstract
In this study, we present seg2med, an advanced medical image synthesis framework that uses Denoising Diffusion Probabilistic Models (DDPM) to generate high-quality synthetic medical images with the conditional anatomical masks from Totalsegmentator. The framework successfully synthesizes CT and MR images from segmentation masks derived from real patient data and XCAT digital phantoms, achieving a Structural Similarity Index Measure (SSIM) of $0.94\pm 0.02$ for CT and $0.89\pm 0.04$ for MR images compared to ground-truth images of real patients, and a Feature Similarity Index Measure (FSIM) of $0.78\pm 0.04$ for CT images from XCAT. The framework demonstrates its generative capability with a Fréchet Inception Distance (FID) of 3.62 for CT image generation. Furthermore, seg2med demonstrates the capability to generate paired multimodal CT and MR images with consistent anatomical structures and performs image conversion between CT and MR modalities, achieving a SSIM of $0.91\pm 0.03$ for the conversion of MR to CT and $0.77\pm 0.04$ for the conversion of CT to MR. Despite the challenges associated with incomplete anatomical information in segmentation masks, the framework achieves promising results, highlighting its potential in cross-modality synthesis and multimodal imaging applications.
The framework also excels at generating high anatomical fidelity CT images, achieving a mean Dice coefficient of $>0.90$ for 11 abdominal organs and $>0.80$ for 34 organs from the 59 organs in 58 test cases, with the highest Dice achieved for the scapula right of $0.96\pm 0.01$. Using the TotalSegmentator toolkit, seg2med enhances its versatility by enabling segmentation mask generation for diverse datasets. This supports applications such as clinical imaging, data augmentation, multimodal image synthesis, and the development of diagnostic algorithms.
@@ I Introduction
Building on the concept of segmentation-guided medical image synthesis, this study presents a framework based on DDPMs. The approach incorporates a preprocessing pipeline that uses TotalSegmentator to generate segmented organ masks, resolves overlapping areas with adjustments, and simulates CT or MRI images for training inputs. The framework evaluates 2-channel DDPMs for multimodal image generation compared to established models such as UNet, CycleGAN, and pix2pix. Comprehensive analyses of the generated images include quantitative metrics, modality conversion efficacy, and their impact on segmentation performance, offering insights into the potential of DDPMs in medical imaging.
@@ II Methods and Materials
II-A Framework
The framework is depicted in Figure 1.
In the Data Preprocessing stage, the TotalSegmentator toolkit is utilized to generate organ segmentation masks from CT and MR images. These masks serve as the foundation for creating paired input data for training. Overlapping regions are resolved using adjustments to ensure accurate representation of anatomical structures. Next, CT Hounsfield Unit (HU) and MR sequence values are allocated to the segmentations, generating CT and MR masks that simulate realistic imaging characteristics. The processed masks are then paired with corresponding CT or MR images, forming the dataset for training the synthesis model.
The Image Synthesis stage employs DDPMs to generate synthetic CT and MR images. The forward diffusion process introduces noise to the input masks, creating progressively noisy versions of the images. The reverse denoising process, guided by a denoising U-Net, reconstructs high-quality synthetic images from these noisy inputs. Separate DDPMs are trained for CT and MR image synthesis to optimize the performance for each modality.
Finally, in the Evaluation stage, the generated images are assessed through comprehensive quantitative and qualitative analyses. Evaluation include segmentation accuracy evaluation and image quality evaluation. The performance of the framework is validated by comparing the synthetic images against ground-truth data, ensuring the effectiveness of the proposed method in multimodal medical imaging tasks.
Figure 1: Illustration for the seg2med framework. The process begins with data preprocessing, including the generation of CT and MR masks. The second stage, image synthesis, employs diffusion probabilistic models (DDPM-CT and DDPM-MR) to generate synthetic MR and CT images. Finally, the evaluation phase assesses the synthetic images using segmentation performance and image quality metrics (e.g., Dice coefficient, PSNR, SSIM, MAE, HistCC, and FSIM). The framework highlights the importance of combining ground-truth-based and non-ground-truth-based evaluations for comprehensive analysis.
II-B Networks
We investigated a conditional denoising diffusion probabilistic model (cDDPM) [19] in this study. It features a diffusion-model-UNet designed with a channel sequence of (64, 128, 256, 256). At each level, two residual blocks were incorporated, and the final layer was enhanced with an attention layer with eight attention heads, each with 32 channels. To address the specific task of image synthesis, the source mask image was appended to the noisy image as an extra channel. Consequently, this modification resulted in a 2-channel input configuration for the UNet component within the diffusion model.
The performance of this 2-channel DDPM on CT image generation (DDPM-CT) was compared with other three supervised generative models: a UNet incorporating attention blocks [20], a pix2pix [11] and a CycleGAN [21]. The UNet, pix2pix, and CycleGAN were trained using their standard settings.
II-C Datasets
@@
In this study, several datasets from different studies were used to add the generalizability of this framework.
@@
The first dataset is the pelvis dataset of task 1 from SynthRAD2023 challenge with a total of 180 patients with paired MR and CT images in the training dataset, and a total of 30 patients with MR images in the validation dataset [22]. The MR images are acquired with a T1 weighted gradient echo sequence. For details on the dataset please refer to [22].
For CT images of 180 patients, the data are divided into 170 training samples and 10 test samples. For MR images of 180 patients, this splitting continues to be used: the images of the same 170 patients are used for training and the rest 10 patients with the 30 patients in the SynthRAD-validation dataset, totally 40 patients, are used for testing.
@@
The third dataset consists of abdomial CT and T1-weighted MRI of the liver. These were acquired on 39 patients as part of the M2OLIE project. This dataset contains 48 CT images and 91 MR T1-weighted Volumetric Interpolated Breath-hold Examination (VIBE) sequence images (in-phase and opposed-phase) in transversal orientation. These images were registered by our previously developed convolutional neural network. For details on the data set please refer to [24]. All the CT images were used for testing. The MRI data of 34 patients (81 image volumes) were used for training and the MRI data of 5 patients (10 image volumes) were used for test. This dataset is referred to as M2OLIE Abdominal CT/MTI in the following.
@@
For the CT framework, by pooling the data, 323 CT image volumes were used for training and 59 CT image volumes for testing. For the MR framework, 251 MR image volumes were used for training and 20 MR image volumes were used for testing. For both frameworks, 56 XCAT phantoms were used for testing only. An overview of the implemented datasets is presented in Table I.
@@
II-D Image Pre-processing
The image pre-processing step involved image normalization, contour regularization, mask generation, and the assignment of simulated grayscale values.
The primary goal of CT image pre-processing was to remove outlier pixels and normalize the images. First, the CT images were windowed using a level of 0 and a width of 2000 to retain lung tissue structures and main bone structures. Subsequently, the images were normalized to a range of [0, 1]. The resolution was resampled to $1.0\times 1.0$ $mm^{2}$ in X-Y plane since this study focused on 2D image generation. For the Z-direction, the resolution was reversed during the stacking of 2D slices to form a 3D volume for each patient, to ensure that the generated slices could be stacked into a volume corresponding to each patient. After windowing and normalization, body contours were extracted from the images to remove irrelevant parts such as arms and the scanner bed, further standardizing the spatial range of the images. The extracted contour map was used not only to regulate the target image but also to ensure consistency in the spatial areas of the generated mask images. Mask generation was a critical step in pre-processing the source images, involving segmentation, value reassignment, and contour regularization. The TotalSegmentator toolkit was employed to generate segmentations of up to 117 anatomical structures from the CT images. Each anatomical class in the segmentation mask was then assigned with an appropriate HU value to create simulated CT masks.
For MR images, pre-processing aimed to normalize intensity values and create accurate segmentation masks. The 10th percentile (p10) and 90th percentile (p90) of intensity values were calculated for each MR volume to define the windowing range. The image intensity values were restricted within this range and then normalized. Similar to the CT images, body contours were extracted and applied to ensure spatial consistency. The TotalSegmentator toolkit was also used for generating MRI segmentation masks. Contour regularization was applied to remove arm regions and ensure alignment with the spatial range of the target images. Next, the VIBE sequence was used to reassign signal intensity values to various organs and tissues within the segmentation map. This step ensured that the simulated MR masks reflected the specific characteristics of the MRI sequence. Consequently, a simulated MR image was generated for each patient.
II-E Experiments
@@
II-E1 Experiment I
First, the four models were trained using 170 pelvis CT images from the SynthRAD training set (SynthRAD Pelvis) [22] and 153 abdomen CT images from the internal dataset (Internal Abdomen CT) [23] for 200 epochs. During training, the input images consisted of CT image masks with reassigned HU values for organ segmentations. The target images were the corresponding ground truth CT images. After training, they were evaluated and compared on 10 CT images of unseen patients from the SynthRAD Pelvis, 10 CT images of unseen patients from Internal Abdomen, and all patients from the M2OLIE Abdomen dataset. The synthetic CT images were assessed by comparison with the ground truth, and the performance of the four implemented models was evaluated specifically in the context of CT image generation. For training the 2-channel DDPM, an initial learning rate of 0.0002, Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and Mean Squared Error (MSE) loss was used.
II-E2 Experiment II
The 2-channel DDPM-MR models were trained using 170 pelvis MR images from the SynthRAD training set (SynthRAD Pelvis) [22] and 81 abdomen MR T1-weighted images from the M2OLIE project (M2OLIE Abdomen) [7]. During early-phase experiments, we observed that the segmentations produced by the total mr task of the Totalsegmentator toolkit were insufficient for synthesizing high-quality MR images. This limitation arose because subcutaneous fat, torso fat, and skeletal muscle are critical for maintaining image fidelity in MR imaging. To address this, we combined the segmentation maps from total mr and tissue types mr tasks into a comprehensive organ map. The merged segmentation maps were then regularized using body contours and assigned MR sequence values generated through the sequence formula. Subsequently, the MR masks were utilized as training inputs, with the corresponding ground truth MR images serving as the target outputs. The training process was conducted over 300 epochs to achieve convergence. Due to differences in acquisition sequences between the two datasets, they were used separately for training and testing. The models were then evaluated on 40 MR images of unseen patient from the SynthRAD Pelvis and 10 MR images of unseen patients from M2OLIE Abdomen.
II-E3 Experiment III
The trained DDPM-CT model was employed to generate CT images using masks from digital phantoms that lacked corresponding ground truth images. CT image generation was performed for 56 Extended Cardiac-Torso (XCAT) digital phantoms [13]. Additionally, a composite digital phantom, createdby merging segmentations and body contour of multiple patients, was used to generate a CT image within our framework, showcasing the model’s robustness and its potential for data augmentation.
II-E4 Experiment IV

### v2

# seg2med: a bridge from artificial anatomy to multimodal medical images

@@ Abstract
We present seg2med, a modular framework for anatomy-driven multimodal medical image synthesis. The system integrates three components to enable high-fidelity, cross-modality generation of CT and MR images based on structured anatomical priors.
First, anatomical maps are independently derived from three sources: real patient data, XCAT digital phantoms, and anatomies-synthetic subjects created by combining organs from multiple patients. Second, we introduce PhysioSynth, a modality-specific simulator that converts anatomical masks into imaging-like prior volumes using tissue-dependent parameters (e.g., HU, T1, T2, $\rho$) and modality-specific signal models. It supports simulation of CT and multiple MR sequences, including GRE, SPACE, and VIBE. Third, the synthesized anatomical priors are used to train 2-channel conditional denoising diffusion probabilistic models (DDPMs), which take the anatomical prior as a structural condition alongside the noisy image, enabling it to generate high-quality, structurally aligned images within its modality.
The framework achieves a Structural Similarity Index Measure (SSIM) of $0.94\pm 0.02$ for CT and $0.89\pm 0.04$ for MR images compared to real patient data, and $0.78\pm 0.04$ FSIM for simulated CT from XCAT. The generative quality is further supported by a Fréchet Inception Distance (FID) of 3.62 for CT synthesis. In modality conversion tasks, seg2med attains SSIM scores of $0.91\pm 0.03$ (MR→CT) and $0.77\pm 0.04$ (CT→MR).
In anatomical fidelity evaluation, synthetic CT images achieve a mean Dice coefficient exceeding 0.90 for 11 key abdominal organs, and over 0.80 for 34 of 59 total organs. These results underscore seg2med’s utility in cross-modality image synthesis, dataset augmentation, and anatomy-aware AI development in medical imaging.
A web-based API demo is available at: https://huggingface.co/spaces/Zeyu0601/frankenstein
@@ I Introduction
In this study, our main contributions are twofold.
(1) We develop an enhanced segmentation preprocessing pipeline that fuses full-organ masks from TotalSegmentator [17] with body contours extracted via OpenCV, producing anatomically faithful segmentation priors that accurately reflect patient-specific anatomy.
(2) We introduce PhysioSynth, a physics-based prior simulation module that converts those segmentation priors into modality-specific physical priors by assigning tissue-level parameters (e.g., HU, T1, T2, $\rho$) and applying MR signal equations. This step goes beyond conventional label-mask inputs, offering fine-grained, interpretable control over modality appearance.
Together, these components enable seg2med to deliver high-fidelity, anatomically aligned cross-modality synthesis with improved flexibility, interpretability, and extensibility.
@@ II Methods and Materials
II-A Datasets
@@
In this study, several datasets from different studies were used to add the generalizability of this framework. In summary, 323 CT image volumes were used for training and 59 CT image volumes for testing, 251 MR image volumes were used for training and 20 MR image volumes were used for testing. 56 XCAT phantoms were used for testing only. An overview of the implemented datasets is presented in Table I.
II-A1 SynthRAD2023
@@
The first dataset is the pelvis dataset of task 1 from SynthRAD2023 challenge with a total of 180 patients with paired MR and CT images in the training dataset, and a total of 30 patients with MR images in the validation dataset [22]. For the MR images, 120 patients in the training set were acquired with a T1-weighted gradient echo (GRE) sequence, and the remaining 60 patients with a T2-weighted SPACE sequence. In the test set, 20 patients were scanned with the T1 GRE sequence and 10 with the T2 SPACE sequence. For further details, please refer to [22].
For CT images of 180 patients, the data aRecently, TotalSegmentator has emerged as a powerful tool for segmenting CT and MRI images,capable of identifying various anatomical structures, including organs, bones, muscles, and vessels. It achieves high accuracy, with a Dice similarity coefficient of 0.943 for CT and 0.824 for MRI, and it adapts to diverse clinical settings while requiring minimal computational resources [17, 18]. Its applications include organ volumetry, disease characterization, radiation therapy planning, and large-scale radiological studies, making it an ideal choice for advancing medical image segmentation workflows. Notably, TotalSegmentator can address the data diversity issue in medical image synthesis by generating unlimited, high-quality anatomical masks from various CT and MR scans. These masks can serve as inputs for synthetic image generation, mitigating the reliance on scarce paired datasets and enabling more robust and generalizable synthesis models. Thus, this study selected TotalSegmentator as the primary segmentation tool for processing real CT and MR images used in this research.
In this study, our main contributions are twofold.
(1) We develop an enhanced segmentation preprocessing pipeline that fuses full-organ masks from TotalSegmentator [17] with body contours extracted via OpenCV, producing anatomically faithful segmentation priors that accurately reflect patient-specific anatomy.
(2) We introduce PhysioSynth, a physics-based prior simulation module that converts those segmentation priors into modality-specific physical priors by assigning tissue-level parameters (e.g., HU, T1, T2, $\rho$) and applying MR signal equations. This step goes beyond conventional label-mask inputs, offering fine-grained, interpretable control over modality appearance.
Together, these components enable seg2med to deliver high-fidelity, anatomically aligned cross-modality synthesis with improved flexibility, interpretability, and extensibility.
III Methods and Materials
III-A Datasets
In this study, several datasets from different studies were used to add the generalizability of this framework. In summary, 323 CT image volumes were used for training and 59 CT image volumes for testing, 251 MR image volumes were used for training and 20 MR image volumes were used for testing. 56 XCAT phantoms were used for testing only. An overview of the implemented datasets is presented in Table I.
III-A1 SynthRAD2023
The first dataset is the pelvis dataset of task 1 from SynthRAD2023 challenge with a total of 180 patients with paired MR and CT images in the training dataset, and a total of 30 patients with MR images in the validation dataset [22]. For the MR images, 120 patients in the training set were acquired with a spoiled T1-weighted gradient echo sequence, and the remaining 60 patients with a T2-weighted SPACE (Sampling Perfection with Application optimized Contrast using different flip angle Evolution) sequence [22]. In the test set, 20 patients were scanned with the T1 GRE sequence and 10 with the T2 SPACE sequence. For further details, please refer togre divided into 170 training samples and 10 test samples. For MR images of 180 patients, this splitting continues to be used: the images of the same 170 patients are used for training and the rest 10 patients with the 30 patients in the SynthRAD-validation dataset, totally 40 patients, are used for testing.
III-A2 Internal Abdominal CT
@@
III-A3 M2OLIE Abdominal CT/MR
The third dataset consists of abdominal CT and T1-weighted MR images acquired from 39 patients as part of the M2OLIE project. It includes 48 CT volumes and 91 T1-weighted MR volumes using Volumetric Interpolated Breath-hold Examination (VIBE) sequences in transversal orientation. The MR data comprises 41 VIBE in-phase, 41 VIBE opposed-phase, 4 VIBE Dixon TRA in-phase, and 5 VIBE Dixon TRA opposed-phase volumes. All images were spatially registered using our previously developed convolutional neural network [24]. Among the MR volumes, 37 in-phase and 37 opposed-phase images were used for training, and 4 in-phase and 4 opposed-phase for testing. Additionally, 3 Dixon in-phase and 4 Dixon opposed-phase volumes were used for training, while 1 Dixon in-phase and 1 Dixon opposed-phase volumes were used for testing. All CT volumes were used exclusively for testing. This dataset is referred to as M2OLIE Abdominal CT/MRI throughout this study.
III-A4 XCAT
@@
III-B Multimodal Data Processing
The framework is depicted in Figure 1.
Figure 1: Illustration for the seg2med framework. The pipeline consists of three modular components. (1) Multimodal Data Processing: Anatomical maps are extracted from real CT/MR images, XCAT phantoms, or seg2med-assembled anatomies using TotalSegmentator and further refined with OpenCV-based body contour detection. (2) PhysioSynth Module: Sequence-specific physical priors are generated based on assigned HU or MR signal equations using tissue-dependent T1, T2, and $\rho$ values. (3) Deep Learning Generation: DDPM-CT and DDPM-MR take the anatomical prior as conditional input to generate structurally aligned, high-fidelity images across CT and multiple MR sequences.
To construct anatomical priors for multimodal image generation, we developed a standardized preprocessing pipeline comprising two steps: (1) full-body segmentation via TotalSegmentator and (2) body contour extraction via OpenCV. These steps were applied separately to CT and MR volumes.
III-B1 Full-Organ and Tissue Segmentation
We employed the TotalSegmentator toolkit to obtain comprehensive anatomical segmentations from CT and MR images. Specifically:
  • For CT images, both the total (full organ) and tissue_types (subcutaneous fat, torso fat, skeletal muscle) modes were used to generate masks covering up to 117 anatomical classes.
  • For MR images, due to differences in contrast and signal characteristics, we used total_mr and tissue_types_mr modes to obtain MR-specific anatomical and tissue segmentations.
These masks were resampled to a consistent resolution of $1.0\times 1.0$ mm2 in the axial plane and stacked to form 3D volumes. Each segmentation was stored as a separate binary map for subsequent simulation and visualization.
III-B2 Body Contour Extraction and Fusion
To isolate the patient’s body region and eliminate irrelevant structures (e.g., scanner bed, arms), we applied contour detection techniques using OpenCV. The process involved:
  1. 1.
    Thresholding the image to isolate the soft-tissue range.
  2. 2.
    Using cv2.findContours() to detect outer edges.
  3. 3.
    Applying cv2.drawContours() and cv2.fillPoly() to generate a solid binary mask of the body contour.
The resulting contour mask was then combined with the anatomical segmentation mask to restrict organ priors to within the physical body region. This fusion ensured consistent spatial alignment across modalities and helped eliminate artifacts from overlapping segmentations. The fused masks served as the input to the PhysioSynth module for further simulation.
III-C PhysioSynth: Physics-Based Prior Simulation
The PhysioSynth module generates modality-specific image priors by converting anatomical segmentation masks into imaging-like volumes using tissue-dependent physical parameters and sequence-specific signal equations.
III-C1 CT Simulation
For CT priors, each organ class in the segmentation mask is assigned a representative Hounsfield Unit (HU) value based on established anatomical attenuation tables. These HU values simulate realistic grayscale intensity distributions for different tissues (e.g., soft tissue, bone, fat, air).
III-C2 MR Simulation
For MRI priors, signal intensities are computed using analytical equations corresponding to the physical behavior of different MR sequences. For each organ, literature-derived T1, T2, and proton density ($\rho$) values are used, and standard sequence parameters (TR, TE, flip angle $\alpha$) are applied. The following equations were used to simulate different T1- and T2-weighted MR sequences:
  • T1-weighted GRE:
    $S=\rho\cdot\sin(\alpha)\cdot\frac{1-e^{-TR/T1}}{1-\cos(\alpha)\cdot e^{-TR/T1}}$ (1)
  • T2-weighted SPACE:
    $S=\rho\cdot e^{-TE/T2}$ (2)
  • T1-weighted VIBE (in-phase and opposed-phase):
    $S=\rho\cdot\sin(\alpha)\cdot\left(\frac{1-e^{-TR/T1}}{1-\cos(\alpha)\cdot e^{-TR/T1}}\right)$ (3)
    Although the same signal equation is used for both in-phase and opposed-phase VIBE images, signal assignments are adjusted for fat–water mixed regions (e.g., liver, subcutaneous fat) to reflect phase interference effects.
  • T1-weighted Dixon VIBE (TRA in-phase and opposed-phase):
    The same GRE-based equation is applied. To simulate the fat–water cancellation in opposed-phase images and enhancement in in-phase images, signal intensity scaling is applied to fat-containing organs.
In all simulations, organ-specific $T1$, $T2$, and $\rho$ values are assigned using lookup tables derived from published studies. The resulting simulated MR priors reflect both anatomical structure and sequence-specific contrast properties, providing a physically meaningful, interpretable prior for training the generative model.
III-D Conditional Diffusion Model with Anatomical Prior
We adopt a conditional denoising diffusion probabilistic model (DDPM) [19] for modality-specific image synthesis. Separate models are trained for CT and MR, each conditioned on anatomical priors derived from segmentation masks.
III-D1 Condition Injection via Mask Concatenation
Let $x_{t}$ denote the noisy target image at time step $t$, and $y$ the anatomical prior image generated from PhysioSynth. The conditional noise prediction function becomes:
$\epsilon_{\theta}(x_{t},t\mid y)$ (4)
To incorporate the anatomical prior, the input to the Diffusion-UNet is constructed by concatenating $x_{t}$ and $y$ as a two-channel input:
$z_{t}=\text{concat}(x_{t},y)\in\mathbb{R}^{2\times H\times W}$ (5)
The time step $t$ is encoded via a sinusoidal positional embedding $\mathbf{e}_{t}$, which is injected into all residual blocks of the Diffusion-UNet, following standard DDPM designs.
III-D2 Training Objective
The training objective is to predict the Gaussian noise $\epsilon$ added to the clean image $x_{0}$, conditioned on the anatomical prior $y$. The simplified DDPM loss function is given by:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{x_{0},\epsilon,t,y}\left[\left\|\epsilon-\epsilon_{\theta}(x_{t},t\mid y)\right\|^{2}\right]$ (6)
where the noisy input is constructed as:
$x_{t}=\sqrt{\bar{\alpha}_{t}}\,x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I)$ (7)
By conditioning on anatomical priors, the model learns to generate structurally consistent medical images aligned with organ-specific spatial context.
III-D3 Training Strategy
To train the DDPM models, we used pairs of anatomical segmentation priors (generated via PhysioSynth) and real images as supervision. During training, a clean target image $x_{0}$ was corrupted using the standard forward diffusion process to obtain the noisy version $x_{t}$, which served as the model input. The anatomical prior was concatenated with $x_{t}$ along the channel dimension, resulting in a two-channel input.
We trained the model for 300 epochs using the Adam optimizer with a learning rate of $2\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. The noise schedule followed a linear $\beta_{t}$ schedule with 500 timesteps. All models were trained using a batch size of 8 on NVIDIA A6000 GPUs.
For comparative analysis, we evaluated the DDPM against three baseline generative models:
  • Attention-UNet [20]: a supervised image-to-image generator enhanced with spatial attention blocks.
  • Pix2Pix [11]: a conditional GAN trained with paired inputs.
  • CycleGAN [21]: an unpaired image translation model adapted to our task using pseudo-label pairs.
All baselines were trained on the same data splits with matched optimizer settings and evaluated using the same set of quantitative metrics (e.g., SSIM, MAE, FID).
III-E Experiments
@@
III-E1 Experiment I
In this experiment, we evaluated the performance of the proposed DDPM model on CT image synthesis (DDPM-CT) and compared it with three baseline generative models: Attention-UNet [20], Pix2Pix [11], and CycleGAN [21].
The models were trained using two CT datasets: 170 pelvis CT images from SynthRAD Pelvis [22] and 153 abdomen CT images from an internal dataset [23]. Anatomical priors were generated as described before.
All models were trained using the same optimizer settings and batch sizes, and evaluated on three testing sets:
  • 10 unseen patients from SynthRAD Pelvis CT
  • 10 unseen patients from Internal Abdomen CT
  • All 48 patients from M2OLIE Abdomen CT
Evaluation metrics included SSIM, MAE, FSIM, and FID scores, comparing synthesized images with real CT ground truth.
III-E2 Experiment II
The second experiment assessed DDPM’s ability to synthesize MR images of varying sequences (DDPM-MR). Training data included:
  • 170 pelvis MR images from SynthRAD Pelvis, including 115 T1-weighted GRE and 55 T2-weighted SPACE sequences
  • 81 T1-weighted MR images from M2OLIE Abdomen, including 37 VIBE in-phase, 37 opposed-phase, 3 Dixon in-phase, and 4 Dixon opposed-phase sequences
The model was tested on:
  • 40 unseen MR images from SynthRAD Pelvis
  • 10 unseen MR images from M2OLIE Abdomen
III-E3 Experiment III
The trained DDPM-CT was employed to generate CT images using masks from digital phantoms that lacked corresponding ground truth images. CT image generation was performed for 56 Extended Cardiac-Torso (XCAT) digital phantoms [13]. Additionally, a composite digital phantom, createdby merging segmentations and body contour of multiple patients, was used to generate a CT image within our framework, showcasing the model’s robustness and its potential for data augmentation.
III-E4 Experiment IV